OpenStudy (gabylovesyou):

@HawkCrimson Factor completely: 6bx + 3ax − 4by − 2ay (3x − 2y) (2b + a) (3x − 2y)(2b − a) (3x + 2y)(2b + a) Prime

OpenStudy (gabylovesyou):

6bx + 3ax : The GCF of 6 and 3 is 3 so it would be 3x

OpenStudy (gabylovesyou):

am i going good?

OpenStudy (gabylovesyou):

lol

OpenStudy (gabylovesyou):

4by - 2ay : The GCF is 2 so it would be 2y

OpenStudy (anonymous):

ur doing great :)

OpenStudy (gabylovesyou):

right now we can take out the third answer... its between A and B

OpenStudy (gabylovesyou):

3x(2b + a) 3 * 2 = 6 so it would be 6bx 3 * a = 3 so it would be 3ax 2y(2b + a) 2 * 2 = 4 so it would be 4by 2 * a = 2 so it would be 2ay

OpenStudy (gabylovesyou):

right?

OpenStudy (anonymous):

-2y(2b + a) besides that ur doing splended :D continue :)

OpenStudy (gabylovesyou):

so then it would be -2y(2b + a) -2 * 2 = -4by -2 * a = -2ay

OpenStudy (gabylovesyou):

is the answer A?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

Do you think you could have solved it if it was 6bx - 2ay + 3ax − 4by ? @Gabylovesyou

OpenStudy (gabylovesyou):

idk dont think so lol